The length of a rectangle is
5 yd longer than its width.
If the perimeter of the rectangle is
34 yd, find its area.

Answer by Alan3354(69443)

The length of a rectangle is
5 yd longer than its width.
If the perimeter of the rectangle is
34 yd, find its area.
----------
Find L & W.
2L + 2W = 34
L = W + 5
---
Area = L*W